OpenStudy (anonymous):

log_4(5x-3)-log_4(2x-1)=3

OpenStudy (anonymous):

log_4(5x-3/2x-1) = 3 5x-3/2x-1 = 4^3 5x-3/2x-1 = 64 5x -3 = 64(2x-1) 5x-3 = 128x-64 128x-5x = 61 123x = 61 x = 61/123

OpenStudy (anonymous):

thanks